# BioCLIP: A Vision Foundation Model for the Tree of Life

### v2

@@ 4.2 Can BioCLIP Generalize to Unseen Taxa?
Taxonomic names are added, removed, and changed as biologists discover and classify new and existing species. BioCLIP should generalize to unseen taxonomic names so that it doesn’t need to be re-trained for every new species. To empirically answer whether BioCLIP generalizes well to unseen taxa, we introduce a new evaluation task that is both biologically and empirically motivated: Rare Species.
@@
Classifying “rare” species is an important and challenging computer vision application in biology, particularly in the context of conservation efforts around the world [79]. To the best of our knowledge, there is no diverse, publicly available rare species classification dataset. Recently published examples [53, 44] lack species diversity with only a dozen classes, compared to our 400. We aim to fill this gap and collect all the species on the IUCN Red List (iucnredlist.org) classified [^5] IUCN has classified $150,388$ species and generally updates their list twice per year (IUCN Update Schedule). The classifications used for this dataset are current as of July 13, 2023. as Near Threatened, Vulnerable, Endangered, Critically Endangered, or Extinct in the Wild. There are approximately $25,000$ species in these categories, and we select $400$ species from this list where each is represented by at least $30$ images in our EOL dataset. These species are then completely removed from TreeOfLife-10M, creating an unseen Rare Species test set with $30$ images per species. This dataset demonstrates (1) BioCLIP’s out-of-distribution generalization to unseen taxa, (2) BioCLIP’s potential applications, and (3) a crucial large Rare Species dataset for the community to address the ongoing biodiversity crisis.
@@
Results. Tab. 4 shows that BioCLIP substantially outperforms both baseline CLIP models, as well as the iNat21-trained CLIP model, at zero-shot classification, especially on unseen taxa (see the column labeled “Rare Species”).
We attribute BioCLIP’s strong zero-shot performance on this broad and diverse set of tasks to the broad and diverse classes present in TreeOfLife-10M. We further explore how data diversity leads to broadly useful image representations in Sec. 4.4. With multimodal contrastive learning (Sec. 3.1) and the incorporation of taxonomic structure (Sec. 3.2), BioCLIP capitalizes on this data diversity for strong zero-shot generalization.
@@
Cross-entropy $16.7$ $26.3$
Hier. cross-entropy $19.3$ $30.7$
CLIP $45.1$ $64.2$
@@
4.3 How Do Text Types Affect Generalization?
We investigate how different text types affect zero-shot generalization by training BioCLIP on a 10% subset of TreeOfLife-10M (10% due to computational constraints). We use our Rare Species dataset because the test classes have every text type, and all species are excluded from training, making it ideal for testing generalization to unseen taxa. Nguyen et al. [54] find that the diversity of captions makes stronger vision models and [69] randomly use one of five different captions for each image during training rather than a single fixed caption. Similarly, we use a mixed text type strategy (Sec. 3.2). How does that affect performance?
Results. The zero-shot ablation results are in Tab. 5; there are several salient observations. First, using taxonomic + common names yields the strongest performance, showing the importance of incorporating the taxonomic structure for generalization. Second, when only using a single text type for training, performance degrades substantially when a different text type is used at test time. Using mixed text types for training, while not the strongest, yields consistently strong performance across the board. These results indicate that mixed text type pre-training largely retains the generalization benefits of using taxonomic names while also providing flexibility of different text types for inference, an important property for a foundation model that may be used for diverse downstream tasks. Finally, using 1M examples from TreeOfLife-10M outperforms using 2.7M examples from iNat21, further confirming the importance of the added data diversity from TreeOfLife-10M.
@@ 4.5 Can BioCLIP Classify More Than Species?
BioCLIP is trained on a (contrastive) species-classification objective, which might limit its downstream use beyond species classification. We consider plant diagnosis with the PlantVillage and PlantDoc datasets, which require classifying both species and the disease present (if any).
Large-scale data labeling is expensive in biology, but biologists always label several instances for use in field guides or museum collections. Few-shot classification is thus an ideal setting for this sort of task transfer.
Results. We find that BioCLIP outperforms baselines at classifying plant diseases based on visual symptoms, in both zero-shot and few-shot settings (see the PlantVillage and PlantDoc columns in Tab. 4).
Notably, while Radford et al. [66] find that CLIP one-shot and two-shot classification is often worse than zero-shot (because few-shot settings cannot use the semantic information in the class name), BioCLIP has learned useful visual representations that are useful even with only one labeled example: BioCLIP’s mean one-shot accuracy is 9.1% higher than zero-shot accuracy.
@@ 4.6 Does BioCLIP Learn the Hierarchy?
BioCLIP demonstrates strong performance in the low-data regime on our extrinsic evaluation, but why? We further conduct an intrinsic evaluation and directly visualize the image feature representations BioCLIP has learned to shed light on this question (Fig. 3). We embed image representations from iNat21’s validation set (which is not see during training) in two dimensions using t-SNE [81], then color the points by the image’s taxonomic label. For each plot, we run t-SNE independently on the subset of examples under the labeled taxonomical rank. Each plot visualizes one rank of the taxonomic hierarchy and the top six categories of the following rank, e.g., the top left plot visualizes the six most common phyla of the Animalia Kingdom. At higher ranks like kingdom (omitted for space) and phylum, both CLIP and BioCLIP have good separation, but BioCLIP’s representation is more fine-grained and contains a richer clustering structure. At lower ranks, BioCLIP produces evidently more separable features, while CLIP’s features are cluttered and lack a clear clustering structure. Appendix F has more qualitative results and visuals.
@@ 5 Related Work
Multimodal foundation model training data. CLIP [66] trained state-of-the-art vision models from noisy, web-scale (100100M+) image-text datasets using a contrastive objective that is optimized for image retrieval. ALIGN [42] and BASIC [62] further scaled the number of training examples from 400M to 6.6B, improving vision representation quality. However, further work [22, 54, 24, 90, 91] all find that dataset diversity and better alignment between the image and caption semantics are more important than dataset size and lead to stronger performance on downstream tasks with less compute. TreeOfLife-10M emphasizes the importance of diversity, adding over 440440K classes to iNat21’s 1010K and leading to stronger zero-shot performance even when training on fewer images (TreeOfLife-1M).
Domain-specific CLIPs. Domain-specific training often beats general training [28, 17]. Labeling domain-specific datasets can be prohibitively expensive because annotators must be subject-matter experts. Image-text training is thus particularly potent because annotations are optional; models can learn from noisy image-text pairs. Ikezogwo et al. [38], Lu et al. [47] gathered 11M+ image-text pairs for use in computational pathology, where expert-labeled examples are difficult to gather due to both time and privacy. We gather $10\times$ the images, emphasizing class diversity.
Hierarchy in computer vision. Hierarchy in computer vision is a well-studied topic, in part because ImageNet [67] concepts are scraped from the hierarchical WordNet [52]. Bilal et al. [10] develop a visual interface to study vision model predictions on ImageNet and find that model confusion patterns follow the same hierarchical structure present in the classes. They leverage this knowledge to incorporate hierarchy into AlexNet’s architecture [43] and improve the top-1 error on ImageNet by 88% absolute. Bertinetto et al. [9] measure image classifiers’ mistakes’ severity and propose alternative training objectives that incorporate class hierarchy. This dramatically reduces mistake severity at the expense of worsening top-1 accuracy. Zhang et al. [93] propose a contrastive learning objective, HiMulCon, where the hierarchical distance between labels corresponds to the desired distance in the embedding space. They use HiMulCon to fine-tune a ResNet that was pre-trained with cross-entropy on ImageNet and outperform cross-entropy on both ImageNet and iNat17 [84].
We repurpose the CLIP objective and feed taxonomic names to an autoregressive text encoder, which forms a dense hierarchical label space into which a vision encoder learns to embed images. This dense label space is better suited to the 454454K unique labels, while prior work applied hierarchies to smaller label spaces.
Computer vision for biology. Fine-grained classification is a classic challenge in computer vision, and biological images are often used to benchmark models. Wah et al. [86], Berg et al. [8], Piosenka [65] all use bird species classification to evaluate fine-grained classification ability. Xiao et al. [89], Cole et al. [19] use biology tasks in contrastive learning frameworks, and Cole et al. [18] uses a biology task in weakly supervised object detection.
@@ 6 Conclusion
Although we leverage the CLIP objective to efficiently learn visual representations over hundreds of thousands of taxa, BioCLIP is fundamentally still trained with a classification objective. In future work, we will further scale up the data, e.g., incorporating research-grade images from inaturalist.org with 100100M+ images, and collect richer textual descriptions of species’ appearances such that BioCLIP can extract fine-grained trait-level representations.
@@ Acknowledgements
The authors would like to thank Josef Uyeda, Jim Balhoff, Dan Rubenstein, Hank Bart, Hilmar Lapp, Sara Beery, and colleagues from the Imageomics Institute and the OSU NLP group for their valuable feedback. We also thank the Bioscan-1M team and the iNaturalist team for making their data available and easy to use, and Jennifer Hammack at EOL for her invaluable help in accessing EOL’s images.
This research was supported by NSF OAC 2118240 and used computing resources from the Ohio Supercomputer Center [15].
@@ Appendix A Reproducibility Statement
We will ensure reproducibility of our results by releasing our datasets (TreeOfLife-10M and Rare Species), data pre-processing code, training code, evaluation code, code to generate all figures (Figs. 2 and 3), and pre-trained model weights. With these resources, anyone with sufficient compute resources can download the original data, then reproduce the pre-processing, training, and evaluation. For those with limited compute, the pre-trained model weights will enable full reproducibility of our evaluation results (Sec. 4).
@@ Appendix E Standard Deviation of Main Results
CLIP $43.7\pm 0.26$ $25.1\pm 0.71$ $21.6\pm 1.05$ $13.7\pm 1.09$ $28.4\pm 0.92$
OpenCLIP $53.7\pm 0.52$ $\mathbf{32.3\pm 0.63}$ $23.2\pm 1.58$ $14.3\pm 0.67$ $29.3\pm 0.68$
BioCLIP $71.8\pm 0.47$ $30.6\pm 0.77$ $\mathbf{57.4\pm 2.4}$ $\mathbf{20.4\pm 1.28}$ $\mathbf{45.3\pm 1.16}$
– iNat21 Only $\mathbf{74.8\pm 0.89}$ $29.6\pm 0.82$ $53.9\pm 0.97$ $19.7\pm 0.80$ $37.6\pm 0.63$
@@
CLIP $73.5\pm 0.37$ $41.2\pm 1.01$ $39.9\pm 0.86$ $24.6\pm 0.90$ $46.9\pm 0.21$
OpenCLIP $81.9\pm 0.25$ $\mathbf{52.5\pm 0.83}$ $42.6\pm 0.82$ $25.0\pm 0.83$ $48.4\pm 0.62$
BioCLIP $90.0\pm 0.12$ $49.3\pm 1.14$ $\mathbf{77.8\pm 0.81}$ $\mathbf{33.6\pm 0.74}$ $\mathbf{66.4\pm 0.16}$
– iNat21 Only $\mathbf{90.1\pm 0.08}$ $48.2\pm 1.24$ $73.7\pm 0.65$ $32.1\pm 1.97$ $56.0\pm 0.36$
@@
BioCLIP $64.5\pm 2.15$ $\mathbf{40.3\pm 3.00}$ $\mathbf{58.8\pm 2.83}$ $\mathbf{84.3\pm 1.90}$ $\mathbf{30.7\pm 1.75}$
@@
BioCLIP $\mathbf{85.6\pm 1.79}$ $\mathbf{62.3\pm 1.82}$ $\mathbf{80.9\pm 1.04}$ $\mathbf{95.9\pm 1.07}$ $\mathbf{47.5\pm 1.35}$
@@ Appendix G More Results of Text-Type
Common $42.4$ $4.4$ $15.8$ $13.3$ $45.2$ $20.7$ $10.7$ $15.4$ $19.6$ $23.1$ $21.0$ 12.0-12.0
Scientific $11.3$ $3.8$ $18.7$ $11.0$ $84.8$ $35.3$ $12.5$ $20.3$ $13.9$ $7.7$ $21.9$ 11.1-11.1
Taxonomic $46.6$ $2.2$ $25.1$ $8.7$ $70.4$ $29.0$ $8.8$ $18.4$ $12.8$ $24.1$ $24.6$ 8.4-8.4
Sci+Com $55.0$ $2.2$ $19.2$ $12.6$ $71.5$ $24.8$ $17.6$ $21.5$ $20.0$ $25.3$ $26.9$ 6.1-6.1
Tax+Com $61.9$ $2.0$ $27.4$ $11.6$ $68.4$ $19.2$ $10.4$ $19.5$ $15.8$ $30.0$ $26.6$ 6.4-6.4
Mixture $66.6$ $3.5$ $30.6$ $17.3$ $86.3$ $32.8$ $19.9$ $18.7$ $24.5$ $30.5$ $33.0$

### v3

@@ 4.2 Can BioCLIP Generalize to Unseen Taxa?
Taxonomic names are added, removed, and changed as biologists discover and classify new and existing species. BioCLIP should generalize to unseen taxonomic names to avoid re-training for every new species. To empirically answer whether BioCLIP generalizes well to unseen taxa, we introduce a new evaluation task that is both biologically and empirically motivated: Rare Species.
@@
Classifying “rare” species is an important and challenging computer vision application in biology, particularly in the context of global conservation efforts [83]. To the best of our knowledge, there is no diverse, publicly available rare species classification dataset. Recently published work [56, 47] lack species diversity with only a dozen classes. We aim to fill this gap, collecting all $\approx 25$K species on the IUCN Red List (iucnredlist.org) classified [^5] IUCN has classified $150,388$ species and generally updates their list twice per year (IUCN Update Schedule). The classifications used for this dataset are current as of July 13, 2023. as Near Threatened, Vulnerable, Endangered, Critically Endangered, or Extinct in the Wild. We select $400$ such species represented by at least $30$ images in our EOL dataset, then remove them from TreeOfLife-10M, creating an unseen Rare Species test set with $30$ images per species. This dataset demonstrates (1) BioCLIP’s out-of-distribution generalization to unseen taxa, (2) BioCLIP’s potential applications, and (3) provides a crucial dataset for the community to address the ongoing biodiversity crisis.
@@
Results. Tab. 4 shows that BioCLIP substantially outperforms both baseline CLIP models, as well as the iNat21-trained CLIP model, at zero-shot classification, especially on unseen taxa (see the “Rare Species” column). We attribute BioCLIP’s strong zero-shot performance on this broad and diverse set of tasks to the broad and diverse classes in TreeOfLife-10M. We explore how data diversity leads to broadly useful image representations in Sec. 4.3.
4.3 How Do Text Types Affect Generalization?
We investigate how different text types affect zero-shot generalization by training BioCLIP on a 10% subset of TreeOfLife-10M (10% due to computational constraints). We use our Rare Species dataset because the test classes have every text type, and all species are excluded from training, making it ideal for testing generalization to unseen taxa. Prior works find that the diversity of captions makes stronger vision models [57] and randomly use one of five different captions for each image during training rather than a single fixed caption [72]. Similarly, we use a mixed text type strategy (Sec. 3.2). How does that affect performance?
Results. The zero-shot ablation results are in Tab. 5; there are several salient observations. First, using taxonomic+common names yields the strongest performance, showing the importance of incorporating the taxonomic structure for generalization. Second, when only using a single text type for training, performance degrades substantially when a different text type is used at test time. Using mixed text types for training yields consistently strong performance across all text types during testing. These results indicate that mixed text type pre-training largely retains the generalization benefits of using taxonomic names while also providing flexibility of different text types for inference, an important property for a foundation model that may be used for diverse downstream tasks. Finally, using 1M examples from TreeOfLife-10M outperforms using 2.7M examples from iNat21, further confirming the importance of the added data diversity from TreeOfLife-10M.
@@ 4.3 How Do Text Types Affect Generalization?
Cross-entropy $16.5$ $26.2$
Hier. cross-entropy $19.3$ $30.5$
CLIP $44.7$ $63.8$
@@ 4.5 Can BioCLIP Classify More Than Species?
BioCLIP is trained on a (contrastive) species-classification objective, which might limit its use beyond species classification. We consider plant diagnosis with the PlantVillage and PlantDoc datasets, which require classifying both species and disease (if any). Large-scale data labeling is expensive, but biologists always label several instances for field guides or museum collections. Few-shot classification is thus an ideal setting for this sort of task transfer.
Results. BioCLIP outperforms baselines at classifying plant diseases based on visual symptoms, in both zero-shot and few-shot settings (see PlantVillage and PlantDoc in Tab. 4). While Radford et al. [69] find that CLIP one-shot and two-shot classification is often worse than zero-shot (because few-shot settings cannot use the semantic information in the class name), BioCLIP has learned useful visual representations that are useful even with only one labeled example: BioCLIP’s mean one-shot accuracy is 9.1% higher than zero-shot accuracy.
@@ 4.6 Does BioCLIP Learn the Hierarchy?
BioCLIP demonstrates strong performance in a low-data regime on our extrinsic evaluation, but why? We further conduct an intrinsic evaluation and visualize BioCLIP’s learned image representations to shed light on this question (Fig. 3). We embed image representations from iNat21’s validation set (unseen during training) using t-SNE [85] and color the points by the image’s taxonomic label. For each plot, we run t-SNE independently on the subset of examples under the labeled taxonomical rank. Each plot visualizes one taxonomic hierarchy rank and the top six categories of the next rank, e.g., the top left plot visualizes the six most common phyla in the Animalia kingdom. At higher ranks like kingdom (omitted for space) and phylum, both CLIP and BioCLIP have good separation, but BioCLIP’s representations are more fine-grained and contain a richer clustering structure. At lower ranks, BioCLIP produces evidently more separable features, while CLIP’s features are cluttered and lack a clear structure. Appendix F has more qualitative results and visuals.
@@ 5 Related Work
Multimodal foundation model training data. CLIP [69] trained state-of-the-art vision models from noisy, web-scale (100100M+) image-text datasets using a contrastive objective that is optimized for image retrieval. ALIGN [45] and BASIC [65] further scaled the number of training examples from 400M to 6.6B, improving vision representation quality. However, further work [24, 57, 26, 93, 94] all find that dataset diversity and better alignment between the image and caption semantics are more important than dataset size and lead to stronger performance on downstream tasks. TreeOfLife-10M emphasizes the importance of diversity, adding over 440440K classes to iNat21’s 1010K and leading to stronger zero-shot performance.
Domain-specific CLIPs. Domain-specific training often beats general training [30, 18], but subject-matter experts are often too expensive to hire to label large-scael domain-specific datasets. Image-text training is thus particularly potent because models can learn from noisy image-text pairs. Ikezogwo et al. [41] and Lu et al. [50] gathered 11M+ image-text pairs for computational pathology. We gather $10\times$ the images, emphasizing class diversity.
Hierarchy in computer vision. Hierarchy in computer vision is well-studied, in part because ImageNet [70] classes are from the hierarchical WordNet [55]. Bilal et al. [10] study model predictions on ImageNet and find that model confusion patterns follow the hierarchical class structure. They incorporate hierarchy into AlexNet’s architecture [46] and improve ImageNet top-1 error by 88% absolute. Bertinetto et al. [9] measure image classifiers’ mistake severity and propose alternative objectives that incorporate hierarchy, reducing mistake severity at the expense of worsening top-1 accuracy. Zhang et al. [96] propose a contrastive objective where the hierarchical distance between labels corresponds to the desired distance in the embedding space, and outperform cross-entropy on ImageNet and iNat17 [88]. We apply hierarchical classification to 454454K unique classes through a repurposed CLIP objective, while prior work applied hierarchies to smaller label spaces.
Computer vision for biology. Fine-grained classification is a classic challenge in computer vision, and biological images are often used to benchmark models. Wah et al. [89], Berg et al. [8], Piosenka [68] all use bird species classification to evaluate fine-grained classification ability. Biology tasks are used for contrastive learning frameworks [92, 20], weakly supervised object detection [19] and semi-supervised learning methods [34].
@@ 6 Conclusion
Although the CLIP objective efficiently learns visual representations over 450450K taxa, BioCLIP is fundamentally trained to do classification. Future work will further scale up the data, e.g., incorporating more than 100100M research-grade images from iNaturalist, and collect richer textual descriptions of species’ appearances such that BioCLIP can extract fine-grained trait-level representations.
@@ Acknowledgements
We thank the Imageomics team (including Josef Uyeda, Jim Balhoff, Dan Rubenstein, Hank Bart, Hilmar Lapp, Sara Beery and Dipanjyoti Paul) and the OSU NLP group for their valuable feedback, the Bioscan-1M and iNaturalist teams for sharing their data, and Jennifer Hammock at EOL for her help accessing EOL’s images. Our research is supported by NSF OAC 2118240 and resources from the Ohio Supercomputer Center [16].
@@
  8. 8.
    Appendix H: Generalized zero-shot learning setting
@@ Appendix A Reproducibility Statement
We ensure reproducibility of our results by releasing our datasets (TreeOfLife-10M and Rare Species), data pre-processing code, training code, evaluation code, code to generate all figures (Figs. 2 and 3), and pre-trained model weights. With these resources, anyone with sufficient compute resources can download the original data, then reproduce the pre-processing, training, and evaluation. For those with limited compute, the pre-trained model weights enable full reproducibility of our evaluation results (Sec. 4).
We provide DOIs as permanent references to our new digital assets:
@@ Appendix C Training Data Aggregation
Despite our efforts, we discovered after training that some hemihomonyms were mislabeled at higher-level taxa (family up to kingdom). This impacts approximately $0.1-0.2\%$ of our data. We are in the process of developing a more robust solution to taxonomic labeling which will also account for re-naming (as is currently in process for many bird species). We intend to release a patch alongside the solution.
@@ Appendix E Standard Deviation of Main Results
CLIP $43.7\pm 0.26$ $25.1\pm 0.71$ $21.6\pm 1.05$ $13.7\pm 1.09$ $28.5\pm 0.65$
OpenCLIP $53.7\pm 0.52$ $32.3\pm 0.63$ $23.2\pm 1.58$ $14.3\pm 0.67$ $29.2\pm 0.64$
Supervised-IN21K $60.2\pm 1.02$ $22.9\pm 0.84$ $14.7\pm 1.38$ $14.4\pm 0.90$ $28.0\pm 0.77$
DINO $40.5\pm 0.96$ $\mathbf{37.0\pm 1.39}$ $23.5\pm 1.49$ $16.4\pm 0.78$ $31.0\pm 0.89$
BioCLIP $71.8\pm 0.47$ $30.6\pm 0.77$ $\mathbf{57.4\pm 2.4}$ $\mathbf{20.4\pm 1.28}$ $\mathbf{44.9\pm 0.73}$
– iNat21 Only $\mathbf{74.8\pm 0.89}$ $29.6\pm 0.82$ $53.9\pm 0.97$ $19.7\pm 0.80$ $36.9\pm 1.02$
@@
CLIP $73.5\pm 0.37$ $41.2\pm 1.01$ $39.9\pm 0.86$ $24.6\pm 0.90$ $46.0\pm 0.33$
OpenCLIP $81.9\pm 0.25$ $52.5\pm 0.83$ $42.6\pm 0.82$ $25.0\pm 0.83$ $47.4\pm 0.34$
Supervised-IN21K $83.9\pm 0.15$ $39.2\pm 1.66$ $32.0\pm 1.90$ $25.4\pm 2.13$ $47.3\pm 0.41$
DINO $70.9\pm 0.34$ $\mathbf{56.9\pm 1.61}$ $46.3\pm 1.37$ $28.6\pm 1.59$ $50.1\pm 0.47$
BioCLIP $90.0\pm 0.12$ $49.3\pm 1.14$ $\mathbf{77.8\pm 0.81}$ $\mathbf{33.6\pm 0.74}$ $\mathbf{65.7\pm 0.43}$
– iNat21 Only $\mathbf{90.1\pm 0.08}$ $48.2\pm 1.24$ $73.7\pm 0.65$ $32.1\pm 1.97$ $55.6\pm 0.16$
@@
Supervised-IN21K $46.7\pm 6.30$ $16.9\pm 2.32$ $\mathbf{62.3\pm 2.28}$ $58.6\pm 4.45$ $27.7\pm 2.86$
DINO $30.7\pm 3.79$ $20.0\pm 1.53$ $60.0\pm 2.15$ $79.2\pm 2.74$ $23.7\pm 2.48$
BioCLIP $64.5\pm 2.15$ $\mathbf{40.3\pm 3.00}$ $58.8\pm 2.83$ $\mathbf{84.3\pm 1.90}$ $\mathbf{30.7\pm 1.75}$
@@
Supervised-IN21K $70.9\pm 2.45$ $30.9\pm 2.64$ $\mathbf{82.4\pm 1.53}$ $82.3\pm 3.81$ $44.7\pm 2.26$
DINO $50.3\pm 3.20$ $34.1\pm 2.87$ $82.1\pm 1.31$ $94.9\pm 1.30$ $40.3\pm 2.32$
BioCLIP $\mathbf{85.6\pm 1.79}$ $\mathbf{62.3\pm 1.82}$ $80.9\pm 1.04$ $\mathbf{95.9\pm 1.07}$ $\mathbf{47.5\pm 1.35}$
@@ Appendix G More Results of Text-Type
Common $58.5$ $4.4$ $15.8$ $13.3$ $45.2$ $20.7$ $10.7$ $15.4$ $19.6$ $24.9$ $22.8$ 10.1-10.1
Scientific $59.7$ $3.8$ $18.7$ $11.0$ $84.8$ $35.3$ $12.5$ $20.3$ $13.9$ $22.3$ $28.2$ 4.7-4.7
Taxonomic $62.7$ $2.2$ $25.1$ $8.7$ $70.4$ $29.0$ $8.8$ $18.4$ $12.8$ $26.6$ $26.4$ 6.5-6.5
Sci+Com $60.2$ $2.2$ $19.2$ $12.6$ $71.5$ $24.8$ $17.6$ $21.5$ $20.0$ $28.0$ $27.7$ 5.2-5.2
Tax+Com $60.2$ $2.0$ $27.4$ $11.6$ $68.4$ $19.2$ $10.4$ $19.5$ $15.8$ $30.4$ $26.4$ 6.5-6.5
Mixture $65.1$ $3.5$ $30.6$ $17.3$ $86.3$ $32.8$ $19.9$ $18.7$ $24.5$ $30.9$ $32.9$